# Hydrodynamics of the electronic Fermi liquid: a pedagogical overview

### v1

@@ I.1 Electron Transport and the Hydrodynamic Regime
$G=N\frac{e^{2}}{h}$ (3)
where $N$ is the number of conducting channels (assuming perfect transmission) and $h$ is Planck’s constant. Due to the long mean free path, ballistic transport is highly non-local; there is no intrinsic meaning to a conductivity per se, as the conductance is manifestly independent of the sample length. This quantum-mechanical result is due to Landauer[13, 11], with a semiclassical generalization due to Sharvin (called Landauer-Sharvin conductance) [14].
@@ I.2 The Hydrodynamic Framework
We take care to distinguish the hydrodynamic regime, as discussed above, from the hydrodynamic framework. The hydrodynamic regime as discussed above amounts to a microscopic specification of the electron-electron scattering rate. By contrast, the hydrodynamic framework is a modeling tool. In the spirit of Landau and Ginzburg, the approach of hydrodynamics is to phenomenologically model a system with a long-wavelength effective theory of (approximately) conserved quantities. This approach is in analogy to the operating principle of thermodynamics, where one uses a handful of macroscopic state variables to summarize and avoid treating an Avogadro’s number of microscopic degrees of freedom. In fact, hydrodynamics can be thought of as a perturbative dynamical extension of equilibrium thermodynamics, with a similarly universal applicability.
@@
where $\nu\equiv\eta/(mn)$ and $\tilde{\zeta}\equiv\zeta/(mn)$ are the kinematic shear and bulk viscosities, respectively. In particular, notice that the Navier-Stokes-Ohm equation [Eq. (10)] is just the momentum equation, or “$F=ma$,” while Eq. (11) is just the continuity equation for the current. Thus, there are only $d+3$ unknowns in this set of equations: the $d$ components of the hydrodynamic velocity $u_{i}$, the particle density $n$, the pressure $p$, and the electric potential $\phi$. In particular, we take the mass $m$, the charge $q$, and the viscosities $\nu$ and $\tilde{\zeta}$ to be given as phenomenological parameters. We note that the system of equations is not yet closed; the state variables $p$ and $\phi$ require their own constitutive relations, which are typically set by equilibrium considerations, i.e. equations of state. With the exception of the ohmic $-\gamma_{\text{mr}}u_{i}$ term, the above equations are structurally identical to those of the hydrodynamics of classical fluids. As we will see, Eq. (10) and Eq. (11) form the framework around which much hydrodynamic phenomena are discussed.
@@ I.3 Why Hydrodynamics Now?
With the concept of hydrodynamics dating back at least to the 1800s, the concept of a hydrodynamic electron fluid is certainly not new. The idea of hydrodynamic electrons in metals is often first associated with Gurzhi’s work in the 1960s [23, 24]. Since then, these ideas have cycled in and out of favor. With the experimental realization of 2D electron gases in the late 1970s, interest in hydrodynamic flow was reignited. Of particular note is Dyakonov and Shur’s work in 1993 [25], where they proposed that the nonlinearities present in the ideal hydrodynamic equations of motion (Euler equations) could lead to the so-called Dyakonov-Shur instability to generate THz waves. In 1995, de Jong and Molenkamp [26] observed hydrodynamic effects in (Al,Ga)As wires. Most recently, a flurry of activity has emerged in the past decade following the discovery of hydrodynamic behavior in a number of ultraclean materials [1, 2, 3]. Monolayer graphene has been the primary material of interest [1, 2, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37], though hydrodynamic signatures have also been reported in other 2DEG systems [26, 38, 39, 40, 41, 42, 43, 44, 45], Weyl semimetals [46, 47, 48], the delafossite PdCoO2 [3, 49, 50], and Sb [51].
@@ II Fermi Liquid Hydrodynamics
[…]
@@
[…]
To obtain dissipative coefficients, we need to keep track of how perturbations $\delta f=f-f_{0}$ decay back to local equilibrium. Formally, such calculations require us to perturbatively solve the (linearized) Boltzmann equation for $\delta f$, which involves explicitly treating $I[f]$ as well as the left-hand side (the “streaming” terms) of Eq. (15). In doing so, one finds corrections to observables due to integration against $f=f_{0}+\delta f$ and therefore to the constitutive relations. For example, computation of the momentum-flux tensor $\Pi_{ij}$ using $f$ leads to an additional correction, namely the the viscous stress tensor $\sigma^{\prime}_{ij}$ (see Eq. (8)). In particular, one finds that $\delta f$ gives the first-order terms responsible for dissipation, e.g. the shear and bulk viscosities. Turning momentum-relaxation back on adds the ohmic dissipation term to the momentum equation. See e.g. Ref. [54] for more details on the Chapman-Enskog approach or Ref. [17] for a modern application to graphene.
@@ II.2 Electrical Resistance
Electrical resistance in a hydrodynamic material is controlled by two dissipative coefficients: the momentum relaxation rate $\gamma_{\text{mr}}$ and the (kinematic shear) viscosity $\nu$ [^5] For time-dependent flows, one also needs to consider the bulk viscosity $\zeta$. In general, viscosity is a rank-4 tensor [see Eq. (36)]. In the linear flow regime, we saw in Sec. I.2 that the Stokes-Ohm equation [Eq. (12)] is equivalent to Ohm’s law with a non-local resistivity $\rho(\mathbf{k})$. While viscous dissipation is fundamentally nonlocal in nature, one can think that in regions of space where the current profile has a spatial gradient, there are two parallel sources of resistivity: one coming from momentum relaxation and another from viscous shear. For example, if we approximate the second derivative on the right-hand side of Eq. (19) by a negative constant $\sim-1/L^{2}$, where $L$ is the typical length scale over which the current varies, then we see that the effect of viscosity is essentially to add a second parallel channel for dissipation in a Matthiessen-rule way: $\gamma_{\text{tot}}\sim\gamma_{\text{mr}}+\nu/L^{2}=\gamma_{\text{mr}}(1+\lambda^{2}/L^{2})$. Physically, one can think that viscous shear adds a second parallel mechanism by which heat is dissipated by the current flow.
@@
Experimentally, anomalous non-ohmic width scaling has been observed in various experiments [3, 46], though we caution that as an effect that depends only on the non-locality of resistivity it need not arise from hydrodynamic electron behavior. These cautionary remarks were particularly pertinent to the hydrodynamic candidate PdCoO2, where experiments had observed anomalous width scaling as evidence of hydrodynamic behavior [3], but upon further inspection concluded that the experiment was likely not in the hydrodynamic regime [73, 50].
@@ II.3 Viscosity
where $u$ is a characteristic velocity and $\ell_{\text{mfp}}$ is a characteristic scattering length (mean free path) [54]. We also remark on the somewhat counterintuitive fact that viscosity diverges as $\ell_{\text{mfp}}\rightarrow\infty$, implying that weakly-interacting gases have very high viscosity. Under the lens that kinematic viscosity is a diffusion constant, this divergence can be rephrased as stating that particles in a weakly-interacting gas travel a long distance before colliding with another particle and diffusing their momentum. In other words, weakly-interacting particles in a medium with no other scattering mechanisms are able to quickly convey their momentum toward boundaries, so that viscous dissipation is more effective  [^6] We assume here the size of the medium $L\gg\ell_{\text{ee}}$ so that the hydrodynamic framework is still valid and viscosity has an unambiguous meaning. See the subsequent discussion on apparent or effective viscosity in this section..
@@ II.4 Superballistic transport
One dramatic signature of non-trivial hydrodynamic behavior is so-called “superballistic transport”. As discussed in Sec. I.1, the conductance in the ballistic regime is given by
$G=N\frac{e^{2}}{h}$ (23)
@@
$G=\frac{ge^{2}}{h}\frac{w}{\lambda_{F}}$ (24)
@@
However, the above results rely on the absence of interactions. In the hydrodynamic regime, one can get around this “fundamental” upper bound; interactions can reduce resistance [97, 103, 104, 105, 106]. This effect was demonstrated most prominently in a constriction or “point-contact” geometry. For ballistic transport, a simple classical picture is that electrons only travel in straight lines. Therefore, only electrons that are aligned with the constriction can pass through, while all others are reflected. For hydrodynamic transport, on the other hand, adjacent layers of fluid are coupled. Fluid passing through the constriction can drag adjacent layers of fluid through the constriction through a viscous force, enabling previously ballistically-forbidden fluid layers to pass through, see Fig. 4. This viscous coupling enhances the conductance of the constriction geometry, allowing for conductances higher than what is allowed by the Landauer-Sharvin formula. This effect is known as superballistic transport. It was experimentally observed in Refs. 28, 33, 34, 37, 107, 108, 109 and serves as a strong signature of interaction-dominated flow.
@@ II.5 The Corbino Geometry and the Corbino Paradox
Figure 5: A plot of current flow and electric potential in the Corbino geometry in the hydrodynamic limit. The expanding flow implies a non-zero viscous stress and therefore non-trivial viscous heating. Despite this production of heat, the electric potential is constant in the bulk of the device (i.e., there is no driving electric field). This apparent inconsistency is known as the Corbino paradox. It is resolved by the sharp drops in voltage at the inner and outer contacts. Adapted from Ref. [111].
@@
$\phi^{c}-\phi=-\frac{2m\nu}{q}K(\hat{\mathbf{n}})(\mathbf{u}\cdot\hat{\mathbf{n}})$ (25)
where $\phi^{c}$ is the contact voltage, $\phi$ is the sample voltage, $m$ is the mass, $\nu$ is the kinematic viscosity, $q$ is the charge, and $K(\hat{\mathbf{n}})$ is the signed extrinsic curvature (inverse radius of the osculating circle) relative to the normal $\hat{\mathbf{n}}$. Qualitatively, one can envision that the viscous stress-energy stored in the boundary curvature within the hydrodynamic sample is converted into a finite voltage in the metallic lead [112]. A key lesson of the Corbino paradox is that one must be much more careful in treating hydrodynamic problems with curved boundaries [113, 112, 102]. The Corbino paradox predictions of a contact voltage drop and the vanishing bulk electric field have been experimentally observed, though much care was taken to isolate only the viscous hydrodynamic component to demonstrate these effects [34].

### v2

@@ I.1 Electron Transport and the Hydrodynamic Regime
$G=N\frac{e^{2}}{2\pi\hbar}$ (3)
where $N$ is the number of conducting channels (assuming perfect transmission). The quantization arises due to the finite width of the sample, analogous to the bound states in the infinite square well. Due to the long mean free path, ballistic transport is highly non-local; there is no intrinsic meaning to a conductivity per se, as the conductance is manifestly independent of the sample length. This quantum-mechanical result is due to Landauer[13, 11], with a semiclassical generalization due to Sharvin (called Landauer-Sharvin conductance) [14].
@@
Approximate momentum conservation in the hydrodynamic regime leads to a host of novel phenomena unusual from the perspective of standard electron transport. For instance, Peierels’ argument implies that a perfectly momentum-conserving hydrodynamic regime would have “infinite conductivity.” It turns out this argument is a bit too quick; though conductivity is ill-defined, hydrodynamic electrons have a separate dissipative mechanism – viscosity – giving rise to a finite conductance in a finite sample with “rough” boundaries. Though somewhat foreign in electron transport, these features are mundane in the classical setting. Fluid flow in free space (or a pipe with smooth boundaries) under a uniform force field is clearly dissipationless, while flow through a pipe with rough boundaries exhibits frictional losses. In this review, we detail some of these phenomena and their distinguishability from traditional ohmic and ballistic transport.
@@ I.2 The Hydrodynamic Framework
We take care to distinguish the hydrodynamic regime, as discussed above, from the hydrodynamic framework. The hydrodynamic regime as discussed above amounts to a microscopic specification of the electron-electron scattering rate. By contrast, the hydrodynamic framework is a modeling tool. In the spirit of Landau and Ginzburg, the approach of hydrodynamics is to phenomenologically model a system with a long-wavelength effective theory of (approximately) conserved quantities. This approach is in analogy to the operating principle of thermodynamics, where one uses a handful of macroscopic state variables to summarize and avoid treating an Avogadro’s number of microscopic degrees of freedom.As a result, the hydrodynamic framework is used in a wide variety of applications beyond particle and charge transport, including spin dynamics, cold atom systems, and integrable models (sometimes under the name “generalized hydrodynamics”) [16, 17, 18]. In fact, hydrodynamics can be thought of as a perturbative dynamical extension of equilibrium thermodynamics, with a similarly universal applicability.
@@
where $\nu\equiv\eta/(mn)$ and $\tilde{\zeta}\equiv\zeta/(mn)$ are the kinematic shear and bulk viscosities, respectively. In particular, notice that the Navier-Stokes-Ohm equation [Eq. (10)] is just the momentum equation, or “F=maF=ma,” while Eq. (11) is just the continuity equation for the current. Thus, there are only $d+3$ unknowns in this set of equations: the $d$ components of the hydrodynamic velocity $u_{i}$, the particle density $n$, the pressure $p$, and the electric potential $\phi$. In particular, we take the mass $m$, the charge $q$, and the viscosities $\nu$ and $\tilde{\zeta}$ to be given as phenomenological parameters. We note that the system of equations is not yet closed; the state variables $p$ and $\phi$ require their own constitutive relations, which are typically set by equilibrium considerations, i.e. equations of state. With the exception of the ohmic $\gamma_{\text{mr}}u_{i}$ term, the above equations are structurally identical to those of the hydrodynamics of classical fluids. As we will see, Eq. (10) and Eq. (11) form the framework around which much hydrodynamic phenomena are discussed.
@@ I.3 Why Hydrodynamics Now?
With the concept of hydrodynamics dating back at least to the 1800s, the concept of a hydrodynamic electron fluid is certainly not new. Hydrodynamics of the electron gas was first studied by Bloch [26] and Jensen [27] in the 1930s, only some 30 years after the discovery of the electron by J.J. Thomson [28]. They emphasized the utility of hydrodynamics to understand the oscillatory “bulk collective response” modes of the interacting electron system, where the fluid formally lived in an infinite, unbounded region. From their work, hydrodynamics was invoked to study, e.g., the various sound modes in liquid He-3 or plasmonic modes in metals [29, 30, 31, 32, 33, 34, 35, 36, 37, 38].
The idea of hydrodynamic electrons in metals as a “confined” fluid, of primary interest for this review, is often first associated with Gurzhi’s work in the 1960s, who emphasized finite-size and boundary effects [39, 40]. Since then, these ideas have cycled in and out of favor. With the experimental realization of 2D electron gases in the late 1970s, interest in hydrodynamic flow was reignited. Of particular note is Dyakonov and Shur’s work in 1993 [41], where they proposed that the nonlinearities present in the ideal hydrodynamic equations of motion (Euler equations) could lead to the so-called Dyakonov-Shur instability to generate THz waves. In 1995, de Jong and Molenkamp [42] observed hydrodynamic effects in (Al,Ga)As wires. Most recently, a flurry of activity has emerged in the past decade following the discovery of hydrodynamic behavior in a number of ultraclean materials [1, 2, 3]. Monolayer graphene has been the primary material of interest [1, 2, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53], though hydrodynamic signatures have also been reported in other 2DEG systems [42, 54, 55, 56, 57, 58, 59, 60, 61], Weyl semimetals [62, 63, 64], the delafossite PdCoO2 [3, 65, 66], and Sb [67].
@@ II Fermi Liquid Hydrodynamics
$\int d\mathbf{k}\,\mathcal{O}(\mathbf{k})\,I[f](\mathbf{k})=0$ (17)
i.e., they vanish upon integration against the collision integral. Integrating the Boltzmann equation [Eq. (15)] multiplied by some collisional invariant $\mathcal{O}$ over momentum $\mathbf{k}$ removes the pesky collision integral. In the absence of fields ($\mathbf{E}=\mathbf{B}=0$), what remains is a continuity equation for the object $\int dk\,\mathcal{O}f\equiv n\langle O\rangle_{f}$, i.e. the density of $\mathcal{O}$ given the distribution $f$. For example, setting $\mathcal{O}=1$ corresponds to $\int d\mathbf{k}\,\mathcal{O}f=n$ the particle density. Thus, integrating both sides of the kinetic equation gives
@@
The constitutive relations can also be derived from the Boltzmann approach. As these require computing observables, doing such derivations requires an explicit solution for the distribution function $f$. While a general solution of the Boltzmann equation is difficult, one can tackle it perturbatively, e.g., using the Chapman-Enskog method [70, 71, 72]. This perturbative expansion is controlled by the Knudsen number $\ell_{\text{mfp}}/L\ll 1$, where $\ell_{\text{mfp}}$ is the characteristic mean free path (including both momentum-conserving and momentum-relaxing collisions) and $L$ is the characteristic sample size.
As a technical remark, in this review we will only concern ourselves with the frequent collision regime $\omega\tau\ll 1$, where $\omega$ is the frequency of the applied field and $\tau\sim v_{F}\ell_{\text{mfp}}$ is a characteristic scattering time  [^4] We remark that since we are working in a finite sample, the limits $\omega\tau\ll 1$ and $\ell_{\text{mfp}}/L\ll 1$ are different. In the study of bulk collective modes, often $L\rightarrow\infty$ so $\omega\tau$ is the only relevant scale of interest. Somewhat confusingly, both $\omega\tau\gg 1$ and $\ell_{\text{mfp}}/L\gg 1$ are called “collisionless” regimes. However, we emphasize that these are physically distinct; in particular, $\ell_{\text{mfp}}/L\gg 1$ requires a careful treatment of boundary scattering, in contrast to $\omega\tau\gg 1$ which can be treated as bulk physics. For perturbative expansions about $\omega\tau\gg 1$ (and discussions regarding the $\omega\tau$ crossover from first to zero zound), see e.g. Ref. [37, 73, 71, 70, 72, 74, 75].. Except when explicitly stated, we take $\omega\rightarrow 0$ and work in steady-state. At present, the cleanest samples have $1/\tau_{\text{mr}}\sim 0.1$ to $1$ THz [2, 21], making it difficult to reach $\omega\tau\gg 1$. Furthermore, most experimentally observed phenomena operate with DC sources. As such, this low-frequency choice is often the experimentally relevant one.
To simplify the discussion, we temporarily ignore momentum relaxation and set $\tau_{\text{mr}}\rightarrow\infty$. The idea of the derivation is as follows. As a zeroth-order ansatz, we guess a distribution function $f_{0}(\mathbf{r},\mathbf{k},t)$ such that the collision integral $I[f_{0}]=0$. At equilibrium where $\mathbf{v}=\mathbf{E}=\mathbf{B}=0$, it is clear that $f_{0}$ is an exact solution. This ansatz is equivalent to assuming local equilibrium; at each point, local collisions alone cannot alter the distribution function. This can be visualized by dividing the fluid into subsystem parcels at the coarse-graining scale, where each subsystem is assumed to establish thermodynamic equilibrium. With $f_{0}$ in hand, one can make explicit calculations of macroscopic observables via integration against $f_{0}$, e.g., $\int dk\mathcal{O}f_{0}$. Therefore, one can use the microscopic definitions of various observables via $f_{0}$ to explicitly derive the constitutive relations; this is analogous to computing equations of state in equilibrium thermodynamics. If one stops at the zeroth-order ansatz $f_{0}$, any relaxational dynamics of collisions are completely ignored since $I[f_{0}]=0$. This results in the Euler equations, i.e. hydrodynamics without dissipation.
To obtain dissipative coefficients, we need to keep track of how perturbations $\delta f=f-f_{0}$ decay back to local equilibrium. Formally, such calculations require us to perturbatively solve the (linearized) Boltzmann equation for $\delta f$, which involves explicitly treating $I[f]$ as well as the left-hand side (the “streaming” terms) of Eq. (15). In doing so, one finds corrections to observables due to integration against $f=f_{0}+\delta f$ and therefore to the constitutive relations. For example, computation of the momentum-flux tensor $\Pi_{ij}$ using $f$ leads to an additional correction, namely the the viscous stress tensor $\sigma^{\prime}_{ij}$ (see Eq. (8)). In particular, one finds that $\delta f$ gives the first-order terms responsible for dissipation, e.g. the shear and bulk viscosities. Turning momentum-relaxation back on adds the ohmic dissipation term to the momentum equation. See, e.g., Ref. [70, 71, 72] for more details on the Chapman-Enskog approach or Ref. [20] for a modern application to graphene.
@@ II.2 Electrical Resistance
Electrical resistance in a hydrodynamic material is controlled by two dissipative coefficients: the momentum relaxation rate $\gamma_{\text{mr}}$ and the (kinematic shear) viscosity $\nu$. [^6] For time-dependent flows, one also needs to consider the bulk viscosity $\zeta$. In general, viscosity is a rank-4 tensor [see Eq. (36)]. In the linear flow regime, we saw in Sec. I.2 that the Stokes-Ohm equation [Eq. (12)] is equivalent to Ohm’s law with a non-local resistivity $\rho(\mathbf{k})$. While viscous dissipation is fundamentally nonlocal in nature, one can think that in regions of space where the current profile has a spatial gradient, there are two parallel sources of resistivity: one coming from momentum relaxation and another from viscous shear. For example, if we approximate the second derivative on the right-hand side of Eq. (19) by a negative constant $\sim-1/L^{2}$, where $L$ is the typical length scale over which the current varies, then we see that the effect of viscosity is essentially to add a second parallel channel for dissipation in a Matthiessen-rule way: $\gamma_{\text{tot}}\sim\gamma_{\text{mr}}+\nu/L^{2}=\gamma_{\text{mr}}(1+\lambda^{2}/L^{2})$. Physically, one can think that viscous shear adds a second parallel mechanism by which heat is dissipated by the current flow.
@@
Experimentally, anomalous non-ohmic width scaling has been observed in various experiments [3, 62], though we caution that as an effect that depends only on the non-locality of resistivity it need not arise from hydrodynamic electron behavior. These cautionary remarks were particularly pertinent to the hydrodynamic candidate PdCoO2, where experiments had observed anomalous width scaling as evidence of hydrodynamic behavior [3], but upon further inspection concluded that the experiment was likely not in the hydrodynamic regime [83, 94, 66].
@@ II.3 Viscosity
where $u$ is a characteristic velocity and $\ell_{\text{mfp}}$ is a characteristic scattering length (mean free path) [70]. We also remark on the somewhat counterintuitive fact that viscosity diverges as $\ell_{\text{mfp}}\rightarrow\infty$ (while keeping the hydrodynamic assumption $\ell_{\text{mfp}}\ll L$), implying that weakly-interacting gases have very high viscosity. Under the lens that kinematic viscosity is a diffusion constant, this divergence can be rephrased as stating that particles in a weakly-interacting gas travel a long distance before colliding with another particle and diffusing their momentum. In other words, weakly-interacting particles in a medium with no other scattering mechanisms are able to quickly convey their momentum toward boundaries, so that viscous dissipation is more effective  [^7] We assume here the size of the medium $L\gg\ell_{\text{ee}}$ so that the hydrodynamic framework is still valid and viscosity has an unambiguous meaning. See the subsequent discussion on apparent or effective viscosity in this section..
@@ II.4 Superballistic transport
One dramatic signature of non-trivial hydrodynamic behavior is so-called “superballistic transport”. As discussed in Sec. I.1, the conductance in the ballistic regime is given by the Landauer formula [Eq. (3)]
$G=N\frac{e^{2}}{2\pi\hbar}$ (23)
@@
$G=\frac{ge^{2}}{2\pi\hbar}\frac{w}{\lambda_{F}}$ (24)
@@
However, the above results rely on the absence of interactions. In the hydrodynamic regime, one can get around this “fundamental” upper bound; interactions can reduce resistance [118, 124, 125, 126, 127]. This effect was demonstrated most prominently in a constriction or “point-contact” geometry. For ballistic transport, a simple classical picture is that electrons only travel in straight lines. Therefore, only electrons that are aligned with the constriction can pass through, while all others are reflected. For hydrodynamic transport, on the other hand, adjacent layers of fluid are coupled. Fluid passing through the constriction can drag adjacent layers of fluid through the constriction through a viscous force, enabling previously ballistically-forbidden fluid layers to pass through, see Fig. 4. The viscous coupling enhances the conductance of the constriction geometry, allowing for conductances higher than what is allowed by the Landauer-Sharvin formula. This effect is known as superballistic transport. It was experimentally observed in Refs. 44, 49, 50, 53, 128, 129, 130; in point-contact geometries, the conductance was observed to exceed the Landauer-Sharvin expectation by more than $15\%$. The observation of superballistic transport serves as a strong signature of interaction-dominated flow.
@@ II.5 The Corbino Geometry and the Corbino Paradox
Figure 5: A plot of current flow and electric potential in the Corbino geometry in the hydrodynamic limit. The expanding flow implies a non-zero viscous stress and therefore non-trivial viscous heating. Despite this production of heat, the electric potential is constant in the bulk of the device (i.e., there is no driving electric field). This apparent inconsistency is known as the Corbino paradox. It is resolved by the sharp drops in voltage at the inner and outer contacts. Adapted from Ref. [132] and Ref. [131].
@@
[…]